# 2-Cats: 2d Copula Approximating Transforms

### v4

@@ 5.1 Empirical First Derivative Estimator
We validate our empirical estimations for the first derivative of Copulas. We use the closed-form equations for the first derivatives of Gaussian, Clayton, and Frank Copulas described in  [57].
These Copulas have a single parameter, which we varied. Gaussian Copulas ($\rho$) has a correlation parameter, and we set it to: 0.1, 0.5, and 0.9. Clayton and Frank copulas’ mixing parameter ($\theta$) was set to 1, 5, and 10. For each of these configurations, we measured the coefficient of determination ($R2$) between empirical estimates and exact formulae. Overall, we found that R2 was above $0.899$ in every setting, validating our approach.
@@
We now turn to our results on synthetic datasets. Our primary focus, as is the case on baseline methods, will be on capturing the data likelihood: $f(x_{1},x_{2})=k_{1}(x_{1})\,k_{2}(x_{2})\,\frac{\partial^{2}H_{\bm{\theta}}(u,v)}{\partial u\partial v}$ Here, $k_{1}(x_{1})$ and $k_{2}(x_{2})$ are KDE models for the PDFs for $X_{1}$ and $X_{2}$ respectivelly. The bandwidth for these methods is determined using Silverman’s rule [59].
@@
Recall that the data density is proportional to the pseudo-likelihood. We primarily emphasized the likelihood loss ($w^{c}$). This accentuates our pivotal component. Notably, the scale of its values differs from that of the MSE components, necessitating a proportional adjustment in weight selection. In light of these factors, we fixed $w^{C}=0.01$, $w^{dC}=0.5$, and $w^{c}=0.1$. We note that these hyper-parameters were sufficient to show how our approach is better than the literature. Thus, no exhaustive grid search was needed. Our models were trained using early stopping. Here, for each training epoch, we evaluated the pseudo-log-likelihood on the training set and kept the weights of the epoch with the best likelihood on the training dataset.
Table 1: Results for Synthetic Data. Copula parameters are shown on the header. Marginals come from two Normal distributions ($\mu=0,\sigma=1$). $95\%$ confidence intervals (bootstrap) show. Green indicates that a 2-Cats model was, on average, better than the best baseline (indicated by Red). Underscores highlight the best statistical ties with the best model. We point out that data is normalized to the [0, 1] range before training. This is a necessity of some deep models such as [38, 51]. Non scaled comparisons are in Appendix E
@@
To assess the performance of our model on synthetic bivariate datasets, we generated Gaussian copulas with correlations $\rho=0.1,0.5,0.9$, and Clayton and Frank copulas with parameters $\theta=1,5,10$. The marginal distributions were chosen as uncorrelated Normal distributions with parameters ($\mu=0,\sigma=1$). We generated 1,500 training samples and 500 test samples. Results are for test samples. Some deep methods require datasets to be scaled to [0, 1] intervals [38, 51]. For this reason, we present results after a min-max scaling of the data.
@@
The winning 2-Cats models are highlighted in the table in Green. The best, on average, baseline methods are in Red. We consider a method better or worse than another when their confidence intervals do not overlap; thus, ties with the best model are shown in underscores. The table shows that 2-Cats is better than baseline methods when the dependency ($\rho$ or $\theta$) increases. For small dependencies, methods not based on Deep Learning outperform Deep ones (including 2-Cats). Nevertheless, 2-Cats is the winning method in 6 out of 9 datasets and is tied with the best in one case.
Table 2: Negative log-likelihood on real datasets. $95\%$ confidence intervals (bootstrap) show. Green indicates that 2-Cats was, on average, better than the best baseline (indicated in Red). Underscores highlight ties with the best model. We point out that data is normalized to the [0, 1] range before training. This is a necessity of some deep models such as [38, 51]
@@
We now turn to real datasets. As was done in previous work, we employ the Boston housing dataset, the Intel-Microsoft (INTC-MSFT), and Google-Facebook (GOOG-FB) stock ratios. These pairs are commonly used in Copula research; in particular, they are the same datasets used by [38, 51]. We employed the same train/test and pre-processing as previous work for fair comparisons.
The results are presented in Table 2, again highlighting the superior performance of the 2-Cats models across all datasets. Only in one setting did the model underperform.
@@ 5.3 On the Lagrangian Term for P3
For the GOOG-FB dataset, significant gains were achieved with the constraints, i.e.: $abs_{u}$ went from $0.17$ to $0.09$ and $abs_{v}$ improved from $0.16$ to $0.03$. $rel_{u}$ improved from $57\%$ to $35\%$ and $rel_{v}$ went from $54\%$ to $14\%$. The negative log-likelihood reduced slightly when Lagrangian was used (from -1.70 to -1.14). Nevertheless, the model is still better than the baselines (see Table 2). Results for the negative log-likelihood do not exactly match the previous ones, as no early stopping was done here.
For INTC-MSOFT, results were similar across $u$ and $v$ dimensions. As in both versions, no significant gains were observed in $abs_{u}$ nor $abs_{v}$; these scores are pretty small (below $0.009$) in both cases. In relative terms, an increase was observed in $abs_{v}$ from roughly 4% to 2%, with $abs_{u}$ being 4% in both cases. No significant changes were observed in the Boston dataset where errors were already minor $abs_{u}$ and $abs_{v}$ below $0.02$ in both instances, with relative errors ranging from 2% to 4% regardless of constraints being employed or not.
These results show that even 2-Cats without constraints may already approximate P3 (Boston and INTC-MSOFT). Constraints may present significant improvements in some cases as GOOG-FB.
@@ 6 Conclusions
In this paper, we presented Copula Approximating Transform models (2-Cats). Different from the literature when training our models; we focus not only on capturing the pseudo-likelihood of Copulas but also on meeting or approximating several other properties, such as the partial derivatives of the $C$ function. Moreover, a second major innovation is proposing estimators that employ using Sobolev training. Overall, our results show the superiority of 2-Cats on synthetic and real datasets.
@@
We point out that before reaching our 2-Cats model, several variations of the model were considered. We refer the reader to Appendix F for results and details of such variations.
@@ Appendix D Conjecture: A Mixture of 2-Cats is a Universal Copula Approximator
Appendix E Non Scaled Synthetic Data
Table 4: Results for non [0, 1] Scaled Data
Gaussian ($\rho$) Clayton ($\theta$) Frank/Joe ($\theta$) 0.1 0.5 0.9 1 5 10 1 5 10 par $2.91\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.09$ $2.65\pm 0.09$ $1.89\pm 0.11$ $1.39\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.10\pm 0.09$ bern $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.09$ $2.66\pm 0.08$ $2.06\pm 0.09$ $1.76\pm 0.08$ $2.78\pm 0.08$ $2.58\pm 0.09$ $2.15\pm 0.08$ pbern $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.10$ $2.67\pm 0.09$ $2.12\pm 0.10$ $1.92\pm 0.08$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.13\pm 0.08$ pspl1 $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.11$ $2.67\pm 0.09$ $2.14\pm 0.18$ $1.61\pm 0.13$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.10\pm 0.09$ pspl2 $2.91\pm 0.09$ $2.69\pm 0.09$ $2.13\pm 0.10$ $2.67\pm 0.09$ $2.02\pm 0.11$ $1.62\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.11\pm 0.09$ TTL0 $2.92\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.09$ $2.65\pm 0.08$ $2.00\pm 0.21$ $1.46\pm 0.09$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.09$ TLL1 $2.92\pm 0.09$ $2.70\pm 0.10$ $2.09\pm 0.09$ $2.65\pm 0.09$ $1.98\pm 0.21$ $1.43\pm 0.10$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.17\pm 0.20$ TLL2 $2.91\pm 0.09$ $2.75\pm 0.20$ $2.08\pm 0.09$ $2.65\pm 0.09$ $2.03\pm 0.25$ $1.41\pm 0.10$ $2.79\pm 0.09$ $2.57\pm 0.09$ $2.11\pm 0.09$ TLL2nn $2.91\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.10$ $2.65\pm 0.09$ $1.93\pm 0.11$ $1.42\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.09$ MR $2.91\pm 0.09$ $2.70\pm 0.09$ $2.16\pm 0.10$ $2.68\pm 0.08$ $2.01\pm 0.11$ $1.54\pm 0.11$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.11\pm 0.08$ probit $2.91\pm 0.09$ $2.69\pm 0.09$ $2.11\pm 0.09$ $2.66\pm 0.08$ $2.05\pm 0.20$ $1.50\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.08$ 2-Cats-L $\mathbf{2.95\pm 0.09}$ $\mathbf{2.79\pm 0.09}$ $1.91\pm 0.10$ $2.67\pm 0.08$ $2.01\pm 0.09$ $0.97\pm 0.10$ $\mathbf{2.90\pm 0.09}$ $\mathbf{2.73\pm 0.10}$ $2.13\pm 0.10$
Table 4 presents results by scaling the input data as is done in Deep Learning methods. The table presents the average negative log-likelihood and the 95% confidence interval. Overall, statistical ties are present throughout the table.
Parametric methods are expected to overperform in this table, as the data was simulated from the model’s dataset. Nevertheless, the statistical ties with such methods show how 2-Cats is able to capture this behavior even if it is a family-free approach. Some cases where 2-Cats outperformed the baselines are shown in bold.
@@ Appendix G Appendix Baseline Source Code
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract discusses the model (Section 1); the properties of the model and how they are met (Section 3 and Appendixes); and empirical results (Section 5 and Appendixes).
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Section 3 discusses how some properties are approximated.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: Section 3 presents our formal proofs. Appendix A complements it.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: Source code is already released in an anonymous link. Different folders present notebooks for our results.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: Source code has already been released. It contains the synthetic data. Real data comes from the literature (also downloaded by our source code).
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Section 5 presents our data splits and random seeds. Our code was also released to verify this. Our model was trained with fixed hyperparameters mentioned in Sections 3 and 5. No tuning was necessary as this was sufficient to overperform baseline methods. The ADAM optimizer is used.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: Our main result tables that compare with baselines present bootstrap confidence intervals.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: Section 5 and Appendix F present how baselines were executed and links to source codes. Our code was executed on Google Colab Pro. Our focus on this work was not on scalability, but we do point out that 2-Cats trains in minutes to hours.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: Our research does not deal with human subjects or sensitive data, nor does it have any copyright issues. We evaluate our research on public non-sensitive and synthetic datasets.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [N/A]
  40. 11.
    Safeguards
  41. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  42. Answer: [N/A]
  43. 12.
    Licenses for existing assets
  44. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  45. Answer: [Yes]
  46. Justification: We only make use of public data and source code for baselines. Our code is already released with am open source license.
  47. 13.
    New Assets
  48. Question: Are new assets introduced in the paper well documented, and is the documentation provided alongside the assets?
  49. Answer: [Yes]
  50. Justification: See the previous question.
  51. 14.
    Crowdsourcing and Research with Human Subjects
  52. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  53. Answer: [N/A]
  54. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  55. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  56. Answer: [N/A]

### v5

@@ 5.1 Empirical First Derivative Estimator
First, we validate our empirical estimations for the first derivative of Copulas. We use the closed-form equations for the first derivatives of Gaussian, Clayton, and Frank Copulas described in  [57].
These Copulas have a single parameter, which we varied. Gaussian Copulas ($\rho$) has a correlation parameter, and we set it to: 0.1, 0.5, and 0.9. Clayton and Frank copulas’ mixing parameter ($\theta$) was set to 1, 5, and 10. For each of these configurations, we measured the coefficient of determination ($R2$) between empirical estimates and exact formulae. Overall, we found that R2 was above $0.899$ in every setting, validating that our estimates are pretty accurate in practive.
@@
We now turn to our main comparisons. Our primary focus, as is the case on baseline methods, will be on capturing the data likelihood (Eq (3)). In our experiments, the PDFs for $X_{1}$ and $X_{2}$ were estimated via KDE. The bandwidth for these methods is determined using Silverman’s rule [59].
@@
Recall that the data density is proportional to the pseudo-likelihood. We primarily emphasized the likelihood loss ($w^{c}$). This accentuates our pivotal component. Notably, the scale of its values differs from that of the MSE components, necessitating a proportional adjustment in weight selection. In light of these factors, we fixed $w^{C}=0.01$, $w^{dC}=0.5$, and $w^{c}=0.1$. We note that these hyper-parameters were sufficient to show how our approach is better than the literature. Our models were trained using early stopping. Here, for each training epoch, we evaluated the pseudo-log-likelihood on the training set and kept the weights of the epoch with the best likelihood on the training dataset.
Table 1: Results for Synthetic Data. Copula parameters are shown on the header. Marginals come from two Normal distributions ($\mu=0,\sigma=1$). $95\%$ confidence intervals (bootstrap) show. Overall, most methods perform statistically the same. This is a positive result; the simulated data comes from well-behaved models. Following the principle of maximum likelihood, it would be unlikely to expect methods to statistically outperform the Parametric (Par) approach.
Gaussian ($\rho$) Clayton ($\theta$) Frank/Joe ($\theta$) 0.1 0.5 0.9 1 5 10 1 5 10 Non-Deep Learn Par $2.91\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.09$ $2.65\pm 0.09$ $1.89\pm 0.11$ $1.39\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.10\pm 0.09$ Bern $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.09$ $2.66\pm 0.08$ $2.06\pm 0.09$ $1.76\pm 0.08$ $2.78\pm 0.08$ $2.58\pm 0.09$ $2.15\pm 0.08$ PBern $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.10$ $2.67\pm 0.09$ $2.12\pm 0.10$ $1.92\pm 0.08$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.13\pm 0.08$ PSPL1 $2.91\pm 0.09$ $2.69\pm 0.09$ $2.15\pm 0.11$ $2.67\pm 0.09$ $2.14\pm 0.18$ $1.61\pm 0.13$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.10\pm 0.09$ PSPL2 $2.91\pm 0.09$ $2.69\pm 0.09$ $2.13\pm 0.10$ $2.67\pm 0.09$ $2.02\pm 0.11$ $1.62\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.11\pm 0.09$ TTL0 $2.92\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.09$ $2.65\pm 0.08$ $2.00\pm 0.21$ $1.46\pm 0.09$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.09$ TLL1 $2.92\pm 0.09$ $2.70\pm 0.10$ $2.09\pm 0.09$ $2.65\pm 0.09$ $1.98\pm 0.21$ $1.43\pm 0.10$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.17\pm 0.20$ TLL2 $2.91\pm 0.09$ $2.75\pm 0.20$ $2.08\pm 0.09$ $2.65\pm 0.09$ $2.03\pm 0.25$ $1.41\pm 0.10$ $2.79\pm 0.09$ $2.57\pm 0.09$ $2.11\pm 0.09$ TLL2nn $2.91\pm 0.09$ $2.69\pm 0.09$ $2.09\pm 0.10$ $2.65\pm 0.09$ $1.93\pm 0.11$ $1.42\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.09$ MR $2.91\pm 0.09$ $2.70\pm 0.09$ $2.16\pm 0.10$ $2.68\pm 0.08$ $2.01\pm 0.11$ $1.54\pm 0.11$ $2.79\pm 0.08$ $2.57\pm 0.09$ $2.11\pm 0.08$ Probit $2.91\pm 0.09$ $2.69\pm 0.09$ $2.11\pm 0.09$ $2.66\pm 0.08$ $2.05\pm 0.20$ $1.50\pm 0.10$ $2.78\pm 0.08$ $2.57\pm 0.09$ $2.12\pm 0.08$ DL NL $1.46\pm 0.08$ $1.32\pm 0.08$ $0.63\pm 0.07$ $1.20\pm 0.06$ $0.47\pm 0.09$ $-0.05\pm 0.10$ $1.39\pm 0.07$ $1.26\pm 0.08$ $0.84\pm 0.09$ IGC $2.92\pm 0.10$ $2.76\pm 0.10$ $2.09\pm 0.10$ $2.64\pm 0.08$ $1.93\pm 0.09$ $1.56\pm 0.11$ $2.87\pm 0.09$ $2.74\pm 0.12$ $2.32\pm 0.13$ Our 2-Cats-L $2.95\pm 0.09$ $2.79\pm 0.09$ $1.91\pm 0.10$ $2.67\pm 0.08$ $2.01\pm 0.09$ $0.97\pm 0.10$ $2.90\pm 0.09$ $2.73\pm 0.10$ $2.13\pm 0.10$ 2-Cats-G $3.07\pm 0.14$ $2.90\pm 0.15$ $1.77\pm 0.10$ $3.04\pm 0.21$ $1.55\pm 0.11$ $0.96\pm 0.10$ $3.13\pm 0.15$ $2.60\pm 0.12$ $2.03\pm 0.13$
@@
To assess the performance of our model on synthetic bivariate datasets, we generated Gaussian copulas with correlations $\rho=0.1,0.5,0.9$, and Clayton and Frank copulas with parameters $\theta=1,5,10$. The marginal distributions were chosen as uncorrelated Normal distributions with parameters ($\mu=0,\sigma=1$). We generated 1,500 training samples and 500 test samples. Results are for test samples. Before continuing, we point out that some deep methods require datasets to be scaled to [0, 1] intervals [38, 51]. Indeed, the author-supplied source code does not execute with data outside of this range (this choice was first used [38, Section 4.2] and followed by  [51, Section 5.1.1]). For fair comparisons, in Appendix E, we present results for models with a min-max scaling of the data.
@@
Initially, note that statistical ties (same average or overlapping intervals) are present throughout the table. The exception is for the NL [13] method, which severely underperforms in this setting, and the methods that did not execute (due to the required data scaling). When reading this table, it is imperative to consider that the simulated data comes from the models evaluated in the Par (first) row. Thus, parametric methods are naturally expected to overperform. Nevertheless, the statistical ties with such methods show how 2-Cats can capture this behavior even if it is a family-free approach.
Table 2: Negative log-likelihood on real datasets. $95\%$ confidence intervals (bootstrap) show. Green indicates that 2-Cats was, on average, better than the best baseline (indicated in Red). Underscores highlight statistical ties with the best model.
@@
We now turn to real datasets. As was done in previous work, we employ the Boston housing dataset, the Intel-Microsoft (INTC-MSFT), and Google-Facebook (GOOG-FB) stock ratios. These pairs are commonly used in Copula research; in particular, they are the same datasets used by [38, 51]. We employed the same train/test and pre-processing as previous work for fair comparisons. Thus, the data is scaled with the same code as the literature, and all deep methods are executed.
The results are presented in Table 2. The winning 2-Cats models are highlighted in the table in Green. The best, on average, baseline methods are in Red. We consider a method better or worse than another when their confidence intervals do not overlap; thus, ties with the best model are shown in underscores. Overall, the table highlights the superior performance of the 2-Cats models across all datasets. Only in one setting, 2-Cats-G on INTC-MSOFT, did the model underperform.
@@ 5.3 On the Lagrangian Term for P3
For the GOOG-FB dataset, significant gains were achieved with the constraints, i.e.: $abs_{u}$ went from $0.17$ to $0.09$ and $abs_{v}$ improved from $0.16$ to $0.03$. $rel_{u}$ improved from $57\%$ to $35\%$ and $rel_{v}$ went from $54\%$ to $14\%$. The negative log-likelihood reduced slightly when Lagrangian was used (from -1.70 to -1.14). Nevertheless, the model is still better than the baselines (see Table 2). Results for the negative log-likelihood do not exactly match the previous ones, as no early stopping was done here. Even though relative errors may appear large, note that such scores are severely sensitive to the tail of the distributions. Here, even a small deviation (e.g., 0.001 to 0.0015) incurs a large relative increase.
For INTC-MSOFT, results were similar across $u$ and $v$ dimensions. As in both versions, no significant gains were observed in $abs_{u}$ nor $abs_{v}$; these scores are pretty small (below $0.009$) in both cases. In relative terms, an increase was observed in $abs_{v}$ from roughly 4% to 2%, with $abs_{u}$ being 4% in both cases. No significant changes were observed in the Boston dataset where errors were already minor $abs_{u}$ and $abs_{v}$ below $0.02$ in both instances, with relative errors ranging from 2% to 4% regardless of the constraints being employed. This shows that 2-Cats may already approximate the Uniform marginals even without constraints.
These results show the efficacy of our Lagrangian penalty. In some settings, such as GOOG-GB, constraints may present significant improvements. Using the Lagrangian term will depend on whether simulation is essential to the end-user (see Appendix B).
@@ 6 Conclusions
In this paper, we presented Copula Approximating Transform models (2-Cats). Different from the literature when training our models; we focus not only on capturing the pseudo-likelihood of Copulas but also on meeting or approximating several other properties, such as the partial derivatives of the $C$ function. Moreover, a second major innovation is proposing Sobolev training for Copulas. Overall, our results show the superiority of 2-Cats on real datasets.
@@ Appendix D Conjecture: A Mixture of 2-Cats is a Universal Copula Approximator
Appendix E Scaled Synthetic Data
Table 4: Results for Synthetic Data. Copula parameters are shown on the header. Marginals come from two Normal distributions ($\mu=0,\sigma=1$). $95\%$ confidence intervals (bootstrap) show. Green indicates that a 2-Cats model was, on average, better than the best baseline (indicated by Red). Underscores highlight the best statistical ties with the best model. We point out that training and test data are independently normalized to the [0, 1] range. This is a necessity of some deep models such as [38, 51] that do not execute on data outside this range (this choice was first used [38, Section 4.2] and followed by  [51, Section 5.1.1]).
Table 4 presents results without scaling the input data as is done in Deep Learning methods. The table presents the average negative log-likelihood and the 95% confidence interval.
The colors and highlights on this table match the main text. The winning 2-Cats models are highlighted in the table in Green. The best, on average, baseline methods are in Red. The table shows that 2-Cats is better than baseline methods when the dependency ($\rho$ or $\theta$) increases. For small dependencies, methods not based on Deep Learning outperform Deep ones (including 2-Cats). Nevertheless, 2-Cats is the winning method in 6 out of 9 datasets and is tied with the best in one case.